# Steady-State Error Compensation for Reinforcement Learning with Quadratic Rewards

### v1

@@ IV-B Reward function
$C_{qua}=w_{1}(\frac{e}{e_{nmax}})^{2}+w_{2}(\frac{u_{r}}{u_{max}})^{2}$ (15)
@@
$C_{abs}=w_{1}\left|\frac{e}{e_{nmax}}\right|+w_{2}\left|\frac{u_{r}}{u_{max}}\right|$ (16)
@@
The PI-reward function is defined as $C_{pi}^{t}=C_{qua}^{t}+C_{I}^{t}$, $C_{I}^{t}=\omega_{4}(\frac{c_{I}^{t}}{c_{nmax}})^{2}$. $C_{I1}^{t+1}$ and $C_{I2}^{t+1}$ correspond to Methods 1 and 2, respectively:
$c_{I1}^{t+1}=\left\{\begin{array}[]{lr}\quad\quad 0\quad\enspace\quad\quad\quad if\quad 0<t\leq t_{threshold}\\ c_{I1}^{t}+\left|e_{t+1}\right|\quad\quad if\quad t>t_{threshold}\end{array}\right.$ (17)
$c_{I2}^{t+1}=c_{I2}^{t}+\kappa\left|e_{t+1}\right|$ (18)
The first term $e_{t+1}$ corresponds to the spacing error in the ACC system and the distance of the vehicle from the centerline of the second lane in the lane-changing system. The second term represents the penalty for actions in both the ACC and lane change systems, while the third term corresponds to the acceleration in the ACC system and the rate of angle change in the lane-changing system. $\omega_{1},\omega_{2},\omega_{3},\omega_{4}$ denotes the weights, and the remaining coefficients are normalization factors. The specific values for all coefficients are provided in the Table I and Table II.
@@
TABLE II: Lane change parameter values
@@
Nominal maximum lateral error $e_{nmax}$ 4m
@@
IV-C Result
Figure 3: Spacing errors in the ACC model
Figure 4: The rate of change in vehicle acceleration for ACC.
First, the results of the ACC model are depicted in Fig. 3 and Fig. 4, where the IPO method illustrates the optimal curve, while reinforcement learning often falls short of achieving such performance. The primary focus is on the final spacing error and acceleration change rate during the tracking process. Utilizing the squared reward function leads to a larger steady-state spacing error, as shown in Fig. 3, and the absolute value reward function induces significant peaks in the acceleration change rate throughout the control process, as shown in Fig. 4. However, employing PI-based methods one and two can reduce the final steady-state spacing error and result in smaller fluctuations in the acceleration change rate.
Figure 5: The distance between the car and the centerline in lane change.
Figure 6: Rate of change in wheel angle for lane change.
For the lane change model results, as illustrated in Fig. 5 and Fig. 6, the focus is on the final distance from the vehicle to the centerline of the second lane and the fluctuation in the wheel angle throughout the entire control process. Comparing the squared reward function with the absolute value reward function, as shown in Fig. 5, there is a difference of five orders of magnitude in the final distance steady-state error, as indicated in Table III. Additionally, the absolute value reward function exhibits larger fluctuations in the wheel angle change rate, as depicted in Fig. 6. The use of PI-based methods 1 and 2 demonstrates superior control performance. Fig. 7 depicts the lane change curves for the six different methods, with specific steady-state error values reflected in Table III.
Figure 7: Lane change curves.
TABLE III: STEADY STATE ERROR
Method ACC Lane change
IPO-qua 4.5e-7m 2.4e-6m
DDPG-qua 2.7e-1m 2.3e-2m
DDPG-pi1 7.3e-3m 1.8e-4m
DDPG-pi2 9.2e-3m 4.9e-4m
IPO-abs 1.1e-9m 4.9e-7m
DDPG-abs 2.9e-3m 5.5e-5m
@@ IV-C Result
This study utilizes PI-based reward functions to improve the control performance of reinforcement learning. The results indicate that using a squared reward function may lead to significant steady-state errors in certain system states, whereas an absolute value reward function may result in an unstable control process with spikes. Adopting PI-based reward functions not only reduces the steady-state error desired but also smoothens the entire control process.

### v2

@@ IV-A2 Reward function
$C_{qua}=w_{1}(\frac{e_{t}}{e_{nmax}})^{2}+w_{2}(\frac{u_{r}}{u_{max}})^{2}$ (13)
@@
$C_{abs}=w_{1}\left|\frac{e_{t}}{e_{nmax}}\right|+w_{2}\left|\frac{u_{r}}{u_{max}}\right|$ (14)
@@
The PI-reward function is defined as $C_{pi}=C_{qua}+C_{I}$, $C_{I}=\omega_{4}(\frac{c_{I}}{c_{nmax}})^{2}$. $C_{I1}$ and $C_{I2}$ correspond to Methods 1 and 2, respectively:
$\mathbf{c}_{I1}=\left\{\begin{array}[]{lr}\enspace\quad\quad\quad\quad 0\quad\quad\quad\quad if\quad 0<t<t_{threshold}\\ \sum_{t_{threshold}}^{t}\left|e_{t}\right|\enspace\quad\quad if\quad t\geq t_{threshold}\end{array}\right.$ (15)
$\mathbf{c}_{I2}=\sum_{0}^{t}\kappa(t)\left|e_{t}\right|$ (16)
The first term $e_{t}$ corresponds to the spacing error in the ACC system. The second term represents the penalty for actions, while the third term corresponds to the acceleration. $\omega_{1},\omega_{2},\omega_{3},\omega_{4}$ denotes the weights, and the remaining coefficients $u_{max}$, $u_{min}$, $e_{nmax}$ and $c_{nmax}$ are normalization factors. The specific values for all coefficients are provided in the Table I.
@@
IV-A3 Result
First, the results of the ACC model are depicted in Fig. 2 and Fig. 3, where the IPO method illustrates the optimal curve, whereas reinforcement learning often falls short of achieving such performance. The primary focus is on the final spacing error and acceleration change rate during the tracking process. Using the quadratic reward function leads to a larger steady-state spacing error, as shown in Fig. 2, and the absolute value reward function induces significant peaks in the acceleration change rate throughout the control process, as shown in Fig. 3. However, employing PI-based Methods 1 and 2 can reduce the final steady-state spacing error and result in small fluctuations in the acceleration change rate. The specific steady-error values and undiscounted episodic rewards (costs) are shown in Table II.
TABLE II: ACC date
Method Steady-state errors [m] Undiscounted episodic
rewards (costs)
IPO-qua 4.5e-7 -9.12
DDPG-qua 2.7e-1 -9.24
DDPG-qua-pi1 7.3e-3 -9.30
DDPG-qua-pi2 9.2e-3 -9.44
IPO-abs 1.1e-9 -19.79
DDPG-abs 2.9e-3 -20.30
Figure 2: Spacing errors in the ACC model. (Only the first 400 timesteps are shown, and the actual training is 600 timesteps)
Figure 3: The rate of change in vehicle acceleration for ACC. (Only the first 400 timesteps are shown, and the actual training is 600 timesteps)
IV-B Lane change control
IV-B1 Model
The vehicle dynamics model used for the lane change in this study is illustrated in Fig. 4.
Figure 4: Dynamic bicycle model for lane change[13].
The state-space equations are as follows:
$\ \dot{x}=u\cos{\varphi}-v\sin{\varphi}$ (17)
$\ \dot{y}=v\cos{\varphi}+u\sin{\varphi}$
$\ \dot{\varphi}=\omega$
$\ \dot{u}=v\omega-\frac{1}{m}F_{Y1}\sin{\delta}$
$\ \dot{v}=-u\omega+\frac{1}{m}(F_{Y1}\cos{\delta}+F_{Y2})$
$\ \dot{\omega}=\frac{1}{I_{z}}(l_{f}F_{Y1}\cos{\delta}-l_{r}F_{Y2})$
$\ \dot{\delta}_{old}=\frac{\delta-\delta_{old}}{\tau}$
Where $x$ and $y$ represent the distances of the vehicle along the $x$ and $y$ axes, $\varphi$ is the angle between the vehicle body and the x-axis, $u$ is the velocity of the vehicle center of mass along the body, $v$ is the velocity of the vehicle center of mass perpendicular to the body, $\omega$ is the angular velocity of the vehicle body, and $\delta_{old}$ stores the decision $\delta$ made by the vehicle in the previous time step. $F_{Y1}$ and $F_{Y2}$ are the lateral forces. Under a mild steering angle, we can assume that:
$\ F_{Y1}\approx k_{f}(\frac{v+l_{f}\omega}{u})-\delta$ (18)
$\ F_{Y2}\approx k_{r}\frac{v-l_{r}\omega}{u}$
The lane change scenario considered in this paper involves the simplest case. The controlled vehicle smoothly approaches the centerline of the second lane. Therefore, the vehicle’s longitudinal acceleration is defined as 0, and the state variable $x$ for longitudinal motion is not considered during the control process.
The distance of the vehicle in the y-direction from the centerline of the second lane is defined as $e$, which serves as a measure of the steady-state error in vehicle control.
For the RL in the lane change environment, the state variables are set as ($e,\varphi,u,v,\omega,\delta_{old}$), and the output actions as $\delta$. The initial state is configured as (4, 0, 30, 0, 0, 0).
IV-B2 Reward function
Similar to the ACC model, only the terms corresponding to nonzero weights in the previous reward function are mentioned. Therefore, the four reward functions are defined as follows:
$C_{qua}=w_{1}(\frac{y_{t}}{y_{nmax}})^{2}+w_{2}(\frac{\delta}{\delta_{max}})^{2}$ (19)
$+w_{3}(\frac{\dot{\delta}_{old}}{(\delta_{max}-\delta_{min})/\tau})^{2}$
$C_{abs}=w_{1}\left|\frac{y_{t}}{y_{nmax}}\right|+w_{2}\left|\frac{\delta}{\delta_{max}}\right|$ (20)
$+w_{3}\left|\frac{\dot{\delta}_{old}}{(\delta_{max}-\delta_{min})/\tau}\right|$
The PI-reward function is defined as $C_{pi}=C_{qua}+C_{I}$, $C_{I}=\omega_{4}(\frac{c_{I}}{c_{nmax}})^{2}$. $C_{I1}$ and $C_{I2}$ correspond to Methods 1 and 2, respectively:
$\mathbf{c}_{I}=\left\{\begin{array}[]{lr}\enspace\quad\quad 0\quad\quad\quad\quad\quad\quad if\quad 0<t<t_{threshold}\\ \sum_{t_{threshold}}^{t}\left|y_{t}\right|\quad\quad\quad if\quad t\geq t_{threshold}\end{array}\right.$ (21)
$\mathbf{c}_{I}=\sum_{0}^{t}\kappa(t)\left|y_{t}\right|$ (22)
The first term $y_{t}$ corresponds to the distance of the vehicle from the centerline of the second lane in the lane-changing system. The second term represents the penalty for actions, and the third term corresponds to the rate of angle change in the lane-changing system. $\omega_{1},\omega_{2},\omega_{3},\omega_{4}$ denotes the weights, and the remaining coefficients $\delta_{max}$, $\delta_{min}$, $y_{nmax}$ and $c_{nmax}$ are normalization factors. The specific values for all coefficients are provided in Table III, and the values of $m,I_{z},k_{f},k_{r},l_{f}$ and $l_{r}$ are referenced in [14].
TABLE III: Lane change parameter values
@@ IV-B2 Reward function
Nominal maximum lateral error $y_{nmax}$ 4m
@@
IV-B3 Result
For the lane change model results, as illustrated in Fig. 5 and Fig. 6, the focus is on the final distance from the vehicle to the centerline of the second lane and the fluctuation in the wheel angle throughout the entire control process. Comparing the quadratic reward function with the absolute value reward function, as shown in Fig. 5, there is a difference of three orders of magnitude in the final distance steady-state error, as indicated in Table IV. In addition, the absolute value reward function exhibits larger fluctuations in the wheel angle change rate, as depicted in Fig. 6. The use of PI-based Methods 1 and 2 with quadratic rewards demonstrate superior control performance. The specific steady-state error values and undiscounted episodic rewards (costs) are reflected in Table IV.
Figure 5: The distance between the car and the centerline in lane change. (Only the first 100 timesteps are shown, and the actual training is 150 timesteps)
Figure 6: Rate of change in wheel angle for lane change. (Only the first 100 timesteps are shown, and the actual training is 150 timesteps)
TABLE IV: Lane change date
Method Steady-state errors [m] Undiscounted episodic
rewards (costs)
IPO-qua 2.4e-6 -3.80
DDPG-qua 2.3e-2 -3.81
DDPG-qua-pi1 1.8e-4 -3.86
DDPG-qua-pi2 4.9e-4 -3.92
IPO-ab 4.5e-7 -6.24
DDPG-abs 5.5e-5 -6.26
@@ IV-B3 Result
This study utilizes PI-based quadratic reward functions by adding an integral term to reduce the steady-state errors of reinforcement learning. The results indicate that using a quadratic reward function may lead to significant steady-state errors in certain systems, whereas an absolute value reward function may result in an unsmooth control process with spikes. Adopting PI-based quadratic reward functions reduces the steady-state errors without inflicting significant spikes of certain system states.